OpenStudy (anonymous):

Analyze the sequence below and complete the instructions that follow. 1, 5, 17, 53, ... Use inductive reasoning to create a general rule for the sequence. A. Multiply by 3, then add 2. B. Multiply by 5. C. Divide by 2, then add 10. D. Multiply by 4, then subtract 2.

OpenStudy (anonymous):

Its asking which one of the answers can you use to get all of the numbers in the sequence For example, lets test B. D says "Multiply by 5" So we if we start with 1 (the first term in the sequence given), and multiply by 5, we get 5, which is the second term of our sequence but if we take 5, and multiply by 5, we get 5*5 = 25, but the third term in the sequence is 17, not 25, so we can see that B wont work for us, because when we plug in 5 into the instructions we are supposed to get 17 So B isn't our answer. your have to test the others and see which one fits

OpenStudy (anonymous):

so it would be A right? 3*1+2= 5 , 3*5+2= 17 , 3*17+2=53

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok thank you ! :)

OpenStudy (anonymous):

You're welcome :)